OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

im guesing my using y=mx+b?

OpenStudy (anonymous):

(y-y1) / (x-x1) = (y2-y1) / (x2-x1) (y-(-4)) / (x-(-1)) = (5-(-4)) / (2-(-4)) (y+4) / (x+1) = (5+4) / (2+4) (y+4) / (x+1) = 9 / 6 (y+4) / (x+1) = 2/3 y+4 = 2(x+1)/3 3(y+4) = 2(x+1) 3y+12 = 2x+2 2x-3y-10=0 or 2x-3y = 10

OpenStudy (anonymous):

donot use that formula use the 2 point form

OpenStudy (anonymous):

i'd used it....

OpenStudy (anonymous):

I am not telling u Romeo

OpenStudy (anonymous):

i was telling pearlruiz

OpenStudy (anonymous):

but is the answer correct? what romeo used?

OpenStudy (anonymous):

yes my answer is correct....

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnk you(: